OpenStudy (asib1214):

How Many grams of sodium chloride at 25C can be dissolved in 150cm^3? Soubility of NaCl = 36g at 25C

OpenStudy (sweetburger):

150cm^3 of water?

OpenStudy (asib1214):

if thats what it sounds like ya sure!!!!

OpenStudy (sweetburger):

well saying its water its usually per 100grams so it should be36g+18g= 54grams in 150cm^3

OpenStudy (asib1214):

actually 150cm^ literally means milliliters.

OpenStudy (sweetburger):

Yes, I understand that.

OpenStudy (sweetburger):

.15 liters sorry so many typos

OpenStudy (asib1214):

aight let me check, i might do my work here and ask u to check it. b bak

OpenStudy (sweetburger):

alright thats fine

OpenStudy (asib1214):

the given solubility is given in 36g/100ml so, if 100ml of water dissolves 36g of NaCl 150ml dissolves=36/100*150=54g